what is the m value for 2px,2py, 2pz orbitals, various type of d orbitals etc.is there any formula to find this?
 

Grade:11

3 Answers

Arun
5 years ago
Orbitals with subshell quantum number l = 1 are called p orbitals. Since the magnetic quantum number m can be -1, 0, or +1 when the value of the subshell quantum number l is one, p orbitals come in sets of three. In each set, one of the orbitals is aligned along each of the three mutually perpendicular axes of the atom; these axes are traditionally designated x, y, and z. The three 2p orbitals are correspondingly designated 2px, 2py, and 2pz

 
Aastha Singh
5 years ago
dear student,
for finding the ‘m’ value, formula is = 2l+1
where ‘l’ is the azimuthal quantum no.
in case of ‘p’ orbital ‘l’ is 1 
therefore 2(1)+1= 3
that means 3 m values including 0 at mean will be= +1, 0, -1
similarly for ‘d’ orbitals, 2l+1 value= 2(2)+ 1= 5
5 m values with 0 at mean = -2, -1, 0 +1, +2
Sadikul Sayed
2 years ago
For 2pz m=0
For 2px m=50% +1,50% -1
For 2py m=50% +1,50% -1
(But I found in a college book mcq part that m=1 for 2py)